OpenStudy (anonymous):

3HBr + Al(OH)3 ----> 3H2O + AlBr3 If 8.56 grams of HBr react with an excess of Al(OH)3 to form 8.91 grams of AlBr3, what is this reaction’s percent yield? Can someone please do it and show me in the simplest and easiest way/steps possible?

OpenStudy (anonymous):

@haseebgul17

OpenStudy (anonymous):

any chance you remember how to do these?

OpenStudy (anonymous):

first find theoretical yield ofAlBr3 molar mass of Al[OH]3=27+[16+1]3=78 molar mass of AlBr3=267 according to equation 78g of Al[OH]3 will give AlBr3=267g 1g of...............................=267/78=3.4g 8.56g of..........................=3.4*8.56=26.8g this is theoretical yield actual yield =8.91g %yield=8.91/26.8*100=33%